# OSDA: A Framework for Open-Set Discovery and Automatic Interpretation of Land-cover in Remote Sensing Imagery

### v1

@@ 1 Introduction
As Earth observation enters the big-data era, the need for intelligent systems that can automatically interpret remote sensing imagery has become pressing  [18]. However, most existing methods adopt a closed-world assumption with a fixed taxonomy, restricting recognition to predefined land-cover classes [11, 64]. This constraint causes substantial performance degradation when models encounter novel objects, rare targets, or dynamically evolving scenes in real-world settings—fundamentally reflecting their limited ability to generalize beyond the trained categories [74].
Consequently, studies within the remote sensing community are increasingly directed toward open-world interpretation. This more challenging yet practically essential task requires models not only to recognize known classes accurately but also to detect, localize, and semantically understand previously unseen targets [57, 59, 75].
@@ 3.1 Dataset Preprocessing
This study employs the OpenEarthMap dataset [63], a global high-resolution land-cover dataset with fine-grained semantic annotations, to fine-tune SAM for the segmentation component. The segmentation phase uses the LoveDA dataset [53] for the test. The LoveDA datset is an urban–rural land-cover segmentation benchmark designed to assess cross-domain generalization. For the first MLLM fine-tuning phase, we use the AID dataset [62], which provides diverse aerial scenes covering a wide range of geographies and imaging conditions. The second fine-tuning stage further incorporates NWPU-RESISC45 [12], a large-scale remote sensing benchmark characterized by substantial intra-class variability; the UC Merced Land-Use dataset [67], a widely used set of high-resolution land-use scenes; PatternNet [77], a benchmark emphasizing discriminative visual patterns for retrieval and classification; and RSI-CB [33], a crowdsourced collection of globally distributed scenes aligned with open geospatial data. Training is further enriched with pixel-level masks generated by SAM [43]. The two-stage fine-tuned MLLM is evaluated on LoveDA dataset [53].
@@ 3.2 Fine-tuning for Segmentation
To support category discovery in the open-set recognition pipeline, we adopt SAM2 [43] as a class‑agnostic segmentation front‑end. We consider two variants: (i) the off‑the‑shelf SAM2 pretrained on large‑scale generic imagery, and (ii) a domain‑adapted SAM2 fine‑tuned on OpenEarthMap [63] using its polygonal land‑cover annotations purely as mask supervision while discarding category names to preclude semantic leakage. Fine‑tuning follows the original SAM2 objective of optimizing promptable, high‑quality region masks [43]. Cross‑domain generalization is assessed on LoveDA [53], where each tile typically contains multiple co‑occurring land‑cover types with intricate boundaries, posing a more demanding scenario than single‑class scenes commonly studied previously. At inference, SAM2 yields pixel‑accurate masks for each region; these structured cues are then consumed by Stage‑II of our MLLM fine‑tuning to guide fine‑grained semantic attribution without exposing any lexical class information, which can improve the Feature Extraction Capabilities of MLLMs. This setup explicitly frames segmentation as an open‑set “unknown‑to‑known” transformation: trained on OpenEarthMap dataset and evaluated on the more topographically complex LoveDA tiles, SAM2 sees no category names yet converts previously unseen content into precise, class‑agnostic region masks that downstream stages can exploit.

### v2

@@ 1 Introduction
As Earth observation enters the big-data era, the need for intelligent systems that can automatically interpret remote sensing imagery has become pressing  [19, 55, 24]. However, most existing methods adopt a closed-world assumption with a fixed taxonomy, restricting recognition to predefined land-cover classes [11, 53, 35]. This constraint causes substantial performance degradation when models encounter novel objects, rare targets, or dynamically evolving scenes in real-world settings—fundamentally reflecting their limited ability to generalize beyond the trained categories [78].
Consequently, studies within the remote sensing community are increasingly directed toward open-world interpretation. This more challenging yet practically essential task requires models not only to recognize known classes accurately but also to detect, localize, and semantically understand previously unseen targets [62, 64, 79, 36, 12].
@@ 3.1 Dataset Preprocessing
This study employs the OpenEarthMap dataset [68], a global high-resolution land-cover dataset with fine-grained semantic annotations, to fine-tune SAM for the segmentation component. The segmentation phase uses the LoveDA dataset [59] for the test. The LoveDA dataset is an urban–rural land-cover segmentation benchmark designed to assess cross-domain generalization. For the first MLLM fine-tuning phase, we use the AID dataset [67], which provides diverse aerial scenes covering a wide range of geographies and imaging conditions. The second fine-tuning stage further incorporates NWPU-RESISC45 [13], a large-scale remote sensing benchmark characterized by substantial intra-class variability; the UC Merced Land-Use dataset [71], a widely used set of high-resolution land-use scenes; PatternNet [81], a benchmark emphasizing discriminative visual patterns for retrieval and classification; and RSI-CB [38], a crowdsourced collection of globally distributed scenes aligned with open geospatial data. Training is further enriched with pixel-level masks generated by SAM [48]. The two-stage fine-tuned MLLM is evaluated on LoveDA dataset [59].
@@ 3.2 Fine-tuning for Segmentation
To support category discovery in the open-set recognition pipeline, we adopt SAM2 [48] as a class‑agnostic segmentation front‑end. We consider two variants: (i) the off‑the‑shelf SAM2 pretrained on large‑scale generic imagery, and (ii) a domain‑adapted SAM2 fine‑tuned on OpenEarthMap [68] using its polygonal land‑cover annotations purely as mask supervision while discarding category names to preclude semantic leakage. Fine‑tuning follows the original SAM2 objective of optimizing promptable, high‑quality region masks [48, 37]. Cross‑domain generalization is assessed on LoveDA [59], where each tile typically contains multiple co‑occurring land‑cover types with intricate boundaries, posing a more demanding scenario than single‑class scenes commonly studied previously. At inference, SAM2 yields pixel‑accurate masks for each region; these structured cues are then consumed by Stage‑II of our MLLM fine‑tuning to guide fine‑grained semantic attribution without exposing any lexical class information, which can improve the Feature Extraction Capabilities of MLLMs. This setup explicitly frames segmentation as an open‑set “unknown‑to‑known” transformation: trained on OpenEarthMap dataset and evaluated on the more topographically complex LoveDA tiles, SAM2 sees no category names yet converts previously unseen content into precise, class‑agnostic region masks that downstream stages can exploit.